Evaluate the following integrals:

5. [ 17+4?) 67 -12 d. sec(Inx) tan(Inx) a. X b. Ja tu ar c. sin(2t)dt

Need all work shown, please and thank you! I will rate!
Answer #1

C Gt-12 J (1+4+-+2) St. 6lt-2 (1+ut-t²3 (2) tyt t² = P differentiate bath sides. - 2t) dit = dp (+-2) dt = -d/ - ()+ - Blue -secl dux). dan (dua) da Let Inces - t sides, differentiate both 1. dx = st = seclt) + c secit tanst.dt . t=ln(s)] = sec (anceSin 12t). at Let et =p differentiate both sides. 2&t= al a) dt = dp When too and t = 37 then poo then p= 317 as sincey Jp - l